Systems Staff Meeting Notes
January 11, 2008
January 11, 2008
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen WilliamsAnnouncements
Review Minutes of 1/4/08
- Awards survey passed around by Debra on behalf of the Employee Recognition Committee.
- Bobbie will be working on her new house this weekend.
- Amy is going to 'Build a Bear' after work. Stephen notes he went to 'Build a Bear' for his 21st birthday (his girlfriend at the time worked there).
- Support Services meeting - security system - they are finishing up putting card swipes in all of the buildings. No more fobs.
- UF closing for the summer semester? It's a possibility.
- Mail will be going to new departmental mailbox numbers. It will be sorted and delivered by UF campus mail.
- The Directors would like a list of all the wonderful things Systems accomplished this year. Please email Will with any suggestions.
- Campus Exchange meeting last week - they are ready to move the first group of adopters over. There was a lot of talk about mailbox restoration. In general, the policy will be that it goes to the dumpster for seven days, where you can restore it yourself. After that, it would have to be restored from tape. Redundancy - they have the hardware and will do it offsite. No cost information as of yet. This was a technical meeting, not one on policy.
- UF emergency contact information - please update it in PeopleSoft.
Ongoing Systems Updates
- Ann #1 - phones - are almost fixed. The message indication light is now working. The rings are close - 4 in front, 3 in the back. Just need them to tweak it to 3 in the front, 3 in the back.
- Ann #2 - Dean Russell is scheduled to attend our next meeting, Friday, Jan. 18.
- Systems training - RIR #2 - did not happen this week.
First group (Will, Debra & Bobbie) to go on Tues 1/15 10AM-12PM. Rm 419 East.
2nd group (Cynthia, Stephen & Chris)
3rd group (Tom B., Winston & Amy)
4th group (Tom M. & Mark)
Groups 2-4 to meet the following consecutive Thursdays, 10AM-12PM.
- Orders are flowing.
- New parts for the old Web server have been received, but are not installed yet.
- Current Library Web presence is now up and running on the virtual server. The old Web server died on Monday morning. We did a restore to VM from an old tape version. On Wednesday, we finished the restores to the current version. Had to install .Net 1.1, recreate shares, the Web cluster reference had to be changed as well as links to: SL2K3Web1.
- McAfee is still running on the old server. Bobbie has the new one installed, but it is not configured yet.
Software Projects Status
- SAN - it is functioning correctly. Logan is working on a speed issue that DLC is experiencing. Might fragmentation be the issue? We have one free license for Disk-Keeper, if needed. Moving files on different roots is excruciatingly slow.
- SMS - Logan did the upgrade. There were some problems with database service accounts. One is using Dewey - would like to move off of that account. Should be using SMS Admin account.
- Windows Updates Slipstreaming - Cynthia is working on the CD. The instructions from Amy helped. It did not recognize her when she logged in as Admin. She will try to log in as herself. She is also trying to reach John Sawyer at CNS for advice, but he has not returned her calls.
- Per Bill Covey: due to budget constraints, we may not be joining Campus AD. We may end up keeping Exchange2K3.
- 2K3SQL - we need to look at CALs. Twelve users are currently connected to it, but we have fifty connections. We are unsure how this works. We need to review all the CALs and licenses and document when they were purchased. We need a better method to track licenses.
- Campus AD public printing demo - Mike from CIRCA is coming on Monday to get the necessary information and square things away. Hopefully, the demo will take place next week. What about the software to track who signs on? It would be like a note - a time stamp.
- Consolidated logging - haven't done too much. Logan is running some tests and will do more soon.
- File server - new way to do user folders. Synching the My Documents folder automatically - on your computer and the server.
- Discussion about possibly having a dummy account to log on to other machines.
- Mark reports an anomaly when logged on as local administrator. If it happens again, inform Will.
DLC Greenstone Issues
- Stephen to take Logan's place in addressing the cross-divisional meeting in discussing blogs and wikis.
- Grover - no updates.
- Tracking Database - the release of the importer should be soon. Mark and Tom are working on it.
- Greenstone - finishing tools before the optimization phase. The Development server should be backed up.
- Circulation Analysis - MSL circulation statistics - worked well, now they want the circ stats for the all of the Libraries. That will be too large to download from FCLA.
- Incident report - Matthew Daley is testing it - haven't heard back.
- Blog - Tom Minton is unable to log into the share. Stephen can log in from his computer, but not from Tom's. Tom needs to be able to access it for the Access Services Web page. Logan may recreate it to see if this helps with the problem.
Students and Department
- SCSI card next week. Also, FTP and monitoring tool hopefully next week.
- Gig switch from CNS for DLC and MSL206 - the request needs to be initiated.
Systems Security Management
- The student schedule is in flux. Should be stabilized by next week.
- Meeting held.